OpenStudy (anonymous):

Central middle school has 2,470 students and the ratio of students to teachers is 26:1. how many new teachers are needed to reduce this ratio to 19:1

OpenStudy (anonymous):

26/1=2470/x

OpenStudy (anonymous):

is this equation right and all i have to do is solve it

OpenStudy (anonymous):

@ganeshie8

OpenStudy (tainted):

2470/26=95 2470/19=130 130-95=35 35 new teachers are needed

OpenStudy (anonymous):

why did you divide it to 19

OpenStudy (tainted):

because you need to reduce the ratio from 26:1 to 19:1 and it asks how many new teachers are needed which is goes from 95 to 130 teachers so it would be 35 new teachers.

OpenStudy (anonymous):

oh i understand thank you so much

OpenStudy (tainted):

No problem!